Point Q is on line segment \overline{PR}

PR


. Given PQ=x,

PQ=x, QR=6,

QR=6, and PR=2x-10,

PR=2x−10, determine the numerical length of \overline{PQ}.

PQ


​.



1
Expert's answer
2021-03-07T17:15:27-0500

Given:

PQ = x

QR = 6

PR = 2x - 10


Since Q is a point on line PR,

"\therefore" PR = PQ + QR

i.e., 2x - 10 = x + 6

2x - x = 6 + 10

"\therefore" x = 16


Hence, the numerical length of PQ = x = 16.